Solve for the specific gravity



1. Coconut oil has density of 930kg/ m³



2. Sea water has a density of 1025 kg/m³

1
Expert's answer
2021-12-20T10:31:54-0500

By the definition of SG, we have:

1)

"SG=\dfrac{\rho_{oil}}{\rho_{w}}=\dfrac{930\ \dfrac{kg}{m^3}}{1000\ \dfrac{kg}{m3}}=0.93"

2)

"SG=\dfrac{\rho_{sea\ water}}{\rho_{w}}=\dfrac{1025\ \dfrac{kg}{m^3}}{1000\ \dfrac{kg}{m3}}=1.025"